I am extremely confused about a problem I'm working on. I can't seem to get the right answer.
A kite flier wondered how high her kite was flying. She used a protractor to measure an angle of 45 degrees from level ground to the kite string. If she used a full 105 yard spool of string, how high, in feet, was the kite? Round the answer to 3 decimal places. (Disregard the string sag and the height of the string reel above the ground.)
I know the correct answer is 222.739 ft, but I don't know how that's possible. Every time I try to solve it I get 24.789ft. Please help!

Answer by nerdybill(7384)

A kite flier wondered how high her kite was flying. She used a protractor to measure an angle of 45 degrees from level ground to the kite string. If she used a full 105 yard spool of string, how high, in feet, was the kite? Round the answer to 3 decimal places. (Disregard the string sag and the height of the string reel above the ground.)
.
You will need to apply trig --
Draw a diagram of the problem, it will help.
You will see that it forms a right triangle with the string as the hypotenuse.
.
Let h = height
.
Since Sin = opposite/hypotenuse
Sin 45 = h/105
h = 105(sin 45)
h = 74.2462 yards
.
Since there are 3 feet per yard:
y = 74.2462*3 = 222.739 feet